# Generic motives and motivic cohomology of fields

### v1

@@ 3.2. Comparison with Beilinson’s approach
The preceding definition can be dualized in an obvious way to define the notion of $t$-injective and derived $t$-injective objects.
To deal with the case of generic motives, one needs to extend slightly the previous definition.
Definition 3.2.7.
[…]
Let $\mathcal{K}$ be a pro-object of $\mathscr{T}$. One says that $\mathcal{K}$ is derived pro-$t$-projective if the following conditions hold:
  1. (i)
    For all $L$ in $\mathscr{T}$, $\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\mathscr{T}}(\mathcal{K},L)\simeq\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\mathscr{T}^{\heartsuit}}(H_{0}(\mathcal{K}),H_{0}(L))$.
  2. (ii)
    The functor $\mathscr{T}^{\heartsuit}\rightarrow\mathscr{A}b,N\mapsto\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\mathscr{T}^{\heartsuit}}(H_{0}(\mathcal{K}),N)$ is exact.
Remark 3.2.8.
The second condition sometimes appears in the literature under the name “pro-projective”.
3.2.9.
Given a stable $\infty$-category $\mathcal{C}$ equipped with a $t$-structure, one can always define a unique $t$-structure on $\operatorname{pro-\mskip-3.0mu}\mathcal{C}$ such that a pro-object $\mathcal{K}$ is positive (resp. negative) if there exists a presentation “$\mathcal{\simeq}\underset{i\in\mathcal{I}}{\text{\rm``lim''}}K_{i}$ such that for all $i\in\mathcal{I}$, $K_{i}$ is positive (resp. negative).
With this definition, $\mathcal{K}$ is derived pro-$t$-projective with respect to $\mathcal{C}$ if and only if the following conditions hold:
  1. (i)
    The pro-object $\mathcal{K}$ is homologically non-$t$-negative in $\operatorname{pro-\mskip-3.0mu}\mathcal{C}$.
  2. (ii)
    For all $L$ in $\mathscr{T}$, $\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\mathscr{T}}(\mathcal{K},L)\simeq\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\mathscr{T}^{\heartsuit}}(\mathcal{K},H_{0}(L))$.
  3. (iii)
    the pro-object $H_{0}(\mathcal{K})$ of $\mathscr{T}^{\heartsuit}$ is projective.
One should be careful however that derived pro-$t$-projective with respect to the $t$-category $\operatorname{Ho}\mathcal{C}$ does not imply derived $t$-projective with respect to the $t$-category $\operatorname{Ho}\operatorname{pro-\mskip-3.0mu}\mathcal{C}$.
Taking into account the previous definition and remark, one gets the following corollary of 3.1.7.
Proposition 3.2.10.
Let $E/k$ be a function field and $i\in\mathbb{Z}$ an integer. Then the generic motive $\operatorname{\hat{M}}(E)\{i\}$ is derived pro-$t$-projective with respect to the homotopy $t$-structure on $\operatorname{DM}(k)$.
Remark 3.2.11.
The dual statement for the cohomological motives $\operatorname{\hat{h}}(E)\{i\}$ in $\operatorname{DM}(k)$ is a bit trickier to state. One has to restrict to geometric motives $M$ for once and use the isomorphism:
$\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{DM_{gm}}(k)}(M,\operatorname{\hat{h}}(E)\{i\})\simeq\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{DM_{gm}}(k)}(\operatorname{\hat{M}}(E)\{-i\},M^{\vee})$
where $M^{\vee}$ denotes the strong dual of $M$. We leave the precise formulation to the reader.
@@ 3.3. Morphisms of generic motives and cycle modules
Generic motives enjoy a rich functorial structure, which parallels that of Milnor K-theory as theorized by Rost in the structure of a cycle module [Ros96, Def 1.1]. In fact, this functoriality can be described by a category ~$\tilde{\mathscr{E}}_{k}$ whose objects are pair $(E,i)$ made by a function field $E$ and an integer $i\in\mathbb{Z}$, and morphisms are described by generators and relations: see the appendix, 6.0.1. Then the aforementioned functoriality can be stated as the existence of a contravariant functor (see op. cit. Th. 5.1.1):
(3.3.1.a) $\rho:(\tilde{\mathscr{E}}_{k})^{\mathrm{op}}\rightarrow\operatorname{Ho}\operatorname{DM^{(0)}}(k),(E,i)\mapsto\operatorname{\hat{M}}(E)\{-i\}.$
Below we will briefly recall the definition of this functor on the generators of morphisms and refer to loc. cit. for the proof of the needed relations. We use notation (D1*), (R1a*) and so on to indicate the data and relations on morphisms of generic motives that correspond to the functor $\rho$, according to the notation of 6.0.1.
@@
Using cohomological motives and given 3.2.4, the above construction amounts to giving a covariant functor:
$\rho^{\vee}:\tilde{\mathscr{E}}_{k}\rightarrow\operatorname{Ho}\operatorname{DM}(k),(E,i)\mapsto\operatorname{\hat{h}}(E)\{i\}.$
The reader might prefer to use this convention (at the price of inverting $e$).
3.3.3.
We start with the first two types of functoriality satisfied by generic motives:
  1. (D1*)
    Covariant functoriality in function fields.– This functoriality simply comes from the fact generic motives are obtained by restriction and $\mathbb{G}_{m}$-twists of the $\infty$-functor $\operatorname{\hat{M}}:\overline{\mathrm{Sm}}_{k}\rightarrow\operatorname{DM_{gm}}(k)$. In particular, an arbitrary morphism $\varphi:E\rightarrow L$ of function fields over $k$ induces a canonical map:
    which actually corresponds to an $\infty$-functor from the category function fields over $k$ to $\operatorname{DM^{(0)}}(k)$.
  2. (D2*)
    Transfers: contravariance in function fields.– Voevodsky’s homological motives are by construction functorial in the category of finite correspondences $\mathrm{Sm}^{\mathrm{cor}}_{k}$. This really corresponds to an $\infty$-functor:
    $\operatorname{M}:\mathrm{Sm}^{\mathrm{cor}}_{k}\rightarrow\operatorname{pro-\mskip-3.0mu}\operatorname{DM_{gm}}(k).$
    One can extend the definition of finite correspondences to essentially smooth $k$-schemes $\mathcal{X}$ and $\mathcal{Y}$. The corresponding abelian group $c(\mathcal{X},\mathcal{Y})$ of finite correspondences from $\mathcal{X}$ to $\mathcal{Y}$ is made by algebraic cycles in $\mathcal{X}\times_{k}\mathcal{Y}$ whose support is finite and equidimensional over $\mathcal{X}$. Composition can be defined as in [Dég07, Def. 1.16] and enjoys the same properties. In particular, one can define a category $\overline{\mathrm{Sm}}^{\mathrm{cor}}_{k}$ of essentially smooth $k$-schemes with theses finite correspondences as morphisms.
    Using the continuity property of [Dég07, Prop. 1.24], one deduces that the preceding functor admits a left Kan extension:
    (3.3.3.a) $\operatorname{\hat{M}}:\overline{\mathrm{Sm}}^{\mathrm{cor}}_{k}\rightarrow\operatorname{pro-\mskip-3.0mu}\operatorname{DM_{gm}}(k).$
    This readily gives the announced functoriality. A finite extension of function field $\varphi:E\rightarrow L$ corresponds to a finite surjective (thus equidimensional) morphism $f:\mathcal{Y}=\operatorname{Spec}(L)\rightarrow\operatorname{Spec}(E)=\mathcal{X}$. Taking the transpose of its graph defines a finite correspondence: ${}^{\mathrm{t}}f\in c(\mathcal{X},\mathcal{Y})$ and therefore a transfer (or Gysin) map:
    Note that this actually gives an $\infty$-functor, from function fields over $k$ equipped with finite morphisms to $\operatorname{DM^{(0)}}(k)$.
Remark 3.3.4.
Note that the functorialities (D1*) and (D2*), together with relations (R1a*), (R1b*), (R1c*) they satisfy, can be gathered using the category $\overline{\mathrm{Sm}}^{\mathrm{cor}}$. In fact, one can consider the full sub-category $\mathrm{Sm}^{\mathrm{cor},(0)}_{k}$ of $\overline{\mathrm{Sm}}^{\mathrm{cor}}$ whose object are the spectrum of function fields over $k$. Then, by restriction of (3.3.3.a), one obtains an $\infty$-functor
$\mathrm{Sm}^{\mathrm{cor},(0)}_{k}\rightarrow\operatorname{DM^{(0)}}(k),\operatorname{Spec}(E)\mapsto\operatorname{\hat{M}}(E)$
which gives a highly structured version of transfers on generic motives. Note moreover that his functors, exactly covers datas (D1*) and (D2*), as well as relations (R1*), of the functor $\rho$.
3.3.5.
Let us now describe the two other kind of functoriality of generic motives:
@@
    (3.3.5.a) $\operatorname{Hom}_{\operatorname{DM^{(0)}}(k)}(\operatorname{\hat{M}}(E),\mathbbm{1}\{n\})\simeq\operatorname{H}_{\mathrm{M}}^{n,n}(E)\simeq K_{n}^{M}(E).$
    Thus, one can associated to any symbol $\sigma=\{u_{1},\ldots,u_{n}\}\in K_{n}^{M}(E)$ a canonical map $\mu(\sigma):\operatorname{\hat{M}}(E)\rightarrow\mathbbm{1}\{n\}$.
    Using the diagonal map $\delta:\operatorname{Spec}(E)\rightarrow\operatorname{Spec}(E)\times_{k}\operatorname{Spec}(E)$, or equivalently the multiplication map $E\otimes_{k}E\rightarrow E$, one deduces the required morphism of generic motives:
    $\gamma_{\sigma}:\operatorname{\hat{M}}(E)\xrightarrow{\delta_{*}}\operatorname{\hat{M}}(E)\otimes\operatorname{\hat{M}}(E)\xrightarrow{\operatorname{Id}\otimes\mu(\sigma)}\operatorname{\hat{M}}(E)\{n\}.$
@@
    Residues.– Let $(E,v)$ be a valued function field over $k$. Let $\mathcal{O}_{v}$ and $\kappa(v)$ be the associated ring of integers and residue (class) field respectively. One deduces a closed immersion between essentially smooth $k$-schemes $i:\mathcal{Z}=\operatorname{Spec}(\kappa(v))\rightarrow\operatorname{Spec}(\mathcal{O}_{v})=\mathcal{X}$, with open complement $j:\mathcal{U}=\operatorname{Spec}(E)\rightarrow\operatorname{Spec}(\mathcal{O}_{v})$. The Gysin triangle extends to this generality and yields a (split) homotopy exact sequence in the stable $\infty$-category $\operatorname{pro-\mskip-3.0mu}\operatorname{DM^{eff}_{gm}}(k)$:
@@
    We call $\partial_{v}$ the residue map associated with $(E,v)$.
This conclude the description of the functor $\rho$ of (3.3.1.a) on morphisms. We refer the reader to [Dég08b] for the proof of the relations.
3.3.6.
Let $\mathbf{E}$ be a motivic complex in $\operatorname{DM}(k)$. Seen as a constant pro-object, it represents a functor $\operatorname{\hat{h}}_{\mathbf{E}}:\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{DM}(k)}(-,\mathbf{E})$, that can be restricted to the category of generic motives. Composing further with the functor $\rho$, one obtains a functor:
$\hat{\mathbf{E}}:\tilde{\mathscr{E}}_{k}\xrightarrow{\rho^{\mathrm{op}}}\operatorname{DM^{(0)}}(k)^{\mathrm{op}}\xrightarrow{\operatorname{\hat{h}}_{\mathbf{E}}}\mathscr{A}b$
which is by definition a cycle premodule. We call it the Rost transform of $\mathbf{E}$. One can check it is in fact a cycle module. Further, according to [Dég11, Th. 3.7], one gets:
@@
$\mathcal{R}:\operatorname{HM}(k)=\operatorname{DM}(k)^{\heartsuit}\rightarrow\operatorname{\mathscr{M}Cycl}_{k},\mathbf{E}\mapsto\hat{\mathbf{E}}$
is an equivalence of categories with quasi-inverse the functor which associates to a cycle module $M$ its 00-th Chow group functor $\underline{H}M:=A^{0}(-,M)$.
@@
Note the last assertion was not proved in loc. cit. but it easily follows as, after inverting $p$, both categories $\operatorname{HM}(k)[1/p]$ and $\operatorname{\mathscr{M}Cycl}_{k}[1/p]$ become invariant under purely inseparable extensions of the base field $k$.
@@
According to 3.1.7, for any motivic complexe $\mathbf{E}$, one gets:
$\hat{\mathbf{E}}(K,n)=\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{DM}(k)}(\operatorname{\hat{M}}(K)\{n\},\mathbf{E})\simeq\operatorname{\underline{H}}_{0,-n}(\mathbf{E})(K).$
In particular, the functor $\mathbf{E}\mapsto\hat{\mathbf{E}}$ factorizes through the homological functor $\operatorname{\underline{H}}_{0,*}$ associated with the homotopy $t$-structure on $\operatorname{DM}(k)$.
Moreover, the abelian group $\hat{\mathbf{E}}(K,n)$ can be interpreted as the fiber of the Nisnevich sheaf $\operatorname{\underline{H}}_{0,-n}(\mathbf{E})$ at the point of the Nisnevich site $\mathrm{Sm}_{k}$ corresponding to the function field $K$. We have interpreted generic motives as points for the $t$-category $\operatorname{DM}(k)$, equipped with its homotopy $t$-structure: 3.1.9, 3.2.10. The Rost transform functor $\mathcal{R}$ is therefore the restriction to this category of points. Remarkably, the restriction a homotopy module $M$ to the category made by these points, enriched via the functoriality defined by the functor $\rho$, allows to fully reconstruct $M$ (as well as morphisms of homotopy modules).
Rost morphisms R* can be interpreted as specialization maps between these various points, but one should be careful they are not isomorphisms. The category $\operatorname{DM^{(0)}}(k)$ is not a groupoid. Nevertheless, from this point of view, cycle modules appear as a generalized version of local systems. Homotopy modules are the sheaf counter-part of these local systems.
3.3.9.
In the PhD of the author ([Dég02, §9.3.2]), the question was raised as to whether morphisms of type (D*) generate all morphisms between generic motives. More precisely, is the above functor $\rho$ full, and/or faithful ?
This hope will be contradicted later (4.1.6), but one can still observe interesting cases where the functor $\rho$ does induce an isomorphism on morphisms. [^22] The next two results was both observed in the PhD of the author [Dég02, §9.3].

### v3

@@ 3.2. Categorical variants and presentations of generic motives
As remarked by the referee, the article [Bei02] uses the opposite category of Voevodsky’s triangulated category $\operatorname{DM_{gm}}(k,\mathbb{Q})$. This can lead to confusion since it is not canonically triangulated. However, the above proposition shows that, due to the good properties of Voevodsky’s (rational) motives, this is not a significant problem. Indeed, in the context of [Bei02], one can work directly in the triangulated category $\operatorname{DM}(k,\mathbb{Q})$, and use the motivic spectrum $\operatorname{\hat{h}}(K)$ in place of the ind-object $\mathcal{C}(\eta)$ in section 3.1 of loc. cit. without altering the essential content of Beilinson’s discussion.
@@ 3.3. Morphisms of generic motives and cycle modules
Generic motives enjoy a rich functorial structure, closely mirroring that of Milnor K-theory as developed by Rost in the framework of cycle modules [Ros96, Def 1.1]. In fact, this functoriality can be described by a category ~$\tilde{\mathcal{E}}_{k}$ whose objects are pairs $(E,i)$ consisting of a function field $E$ and an integer $i\in\mathbb{Z}$, and morphisms are described by generators and relations: see the appendix, 7.0.1. Then the aforementioned functoriality can be expressed through the existence of a contravariant functor (see [Dég08b, Th. 5.1.1]):
(3.3.1.a) $\rho:(\tilde{\mathcal{E}}_{k})^{\mathrm{op}}\rightarrow\operatorname{Ho}\operatorname{\mathscr{D}M^{(0)}}(k),(E,i)\mapsto\operatorname{\hat{M}}(E)\{-i\}.$
We will use this functor in the next section so we briefly recall its definition on the generators of the morphisms, and refer to loc. cit. for the proof of the needed relations. We use notation (D1*), (R1a*), and so on, to indicate the data and relations on morphisms of generic motives that correspond to the functor $\rho$, according to the notation of 7.0.1.
3.3.2.
We start with the first two types of functorialities satisfied by generic motives:
  1. (D1*)
    Covariant functoriality in function fields.– This functoriality simply arises from the fact that generic motives are obtained by restriction and $\mathbb{G}_{m}$-twists of the $\infty$-functor $\operatorname{\hat{M}}:\overline{\mathrm{Sm}}_{k}\rightarrow\operatorname{\mathscr{D}M_{gm}}(k)$. In particular, an arbitrary morphism $\varphi:E\rightarrow L$ of function fields over $k$ induces a canonical map:
    which actually defines an $\infty$-functor from the category function fields over $k$ to $\operatorname{\mathscr{D}M^{(0)}}(k)$.
  2. (D2*)
    Transfers: contravariance in function fields.– By construction, Voevodsky’s homological motives are functorial in the category of finite correspondences $\mathrm{Sm}^{\mathrm{cor}}_{k}$. This gives rise to an $\infty$-functor:
    $\operatorname{M}:\mathrm{Sm}^{\mathrm{cor}}_{k}\rightarrow\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M_{gm}}(k).$
    The definition of finite correspondences can be extended to essentially smooth $k$-schemes $\mathcal{X}$ and $\mathcal{Y}$. The corresponding abelian group $c(\mathcal{X},\mathcal{Y})$ of finite correspondences from $\mathcal{X}$ to $\mathcal{Y}$ is composed of algebraic cycles in $\mathcal{X}\times_{k}\mathcal{Y}$ whose support is finite and equidimensional over $\mathcal{X}$. Composition is defined as in [Dég07, Def. 1.16] and has the same properties. Thus, one can define a category $\overline{\mathrm{Sm}}^{\mathrm{cor}}_{k}$ of essentially smooth $k$-schemes with finite correspondences as morphisms.
    Using the continuity property of [Dég07, Prop. 1.24], one deduces that the preceding functor admits a right Kan extension:
    (3.3.2.a) $\operatorname{\hat{M}}:\overline{\mathrm{Sm}}^{\mathrm{cor}}_{k}\rightarrow\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M_{gm}}(k).$
    This immediately provides the desired functoriality. A finite extension of function fields $\varphi:E\rightarrow L$ corresponds to a finite surjective (and hence equidimensional) morphism $f:\mathcal{Y}=\operatorname{Spec}(L)\rightarrow\operatorname{Spec}(E)=\mathcal{X}$. Taking the transpose of its graph defines a finite correspondence ${}^{\mathrm{t}}f\in c(\mathcal{X},\mathcal{Y})$ and therefore a transfer (or Gysin) map:
    Note that this construction defines an $\infty$-functor, from the category of function fields over $k$, equipped with finite morphisms to $\operatorname{\mathscr{D}M^{(0)}}(k)$.
@@
Note that the functorial properties (D1*) and (D2*), together with the relations (R1a*), (R1b*), (R1c*) they satisfy, can be obtained using the category $\overline{\mathrm{Sm}}^{\mathrm{cor}}$. In fact, one can consider the full subcategory $\mathrm{Sm}^{\mathrm{cor},(0)}_{k}$ of $\overline{\mathrm{Sm}}^{\mathrm{cor}}$ whose objects are the spectra of function fields over $k$. Then, by restricting (3.3.2.a), one obtains an $\infty$-functor
$\mathrm{Sm}^{\mathrm{cor},(0)}_{k}\rightarrow\operatorname{\mathscr{D}M^{(0)}}(k),\operatorname{Spec}(E)\mapsto\operatorname{\hat{M}}(E)$
which gives a highly structured version of transfers on generic motives. Moreover, note that this functor precisely recovers the data (D1*) and (D2*), as well as the relations (R1*), of the functor $\rho$.
3.3.4.
Let us now describe the two other types of functoriality for generic motives:
@@
    (3.3.4.a) $\operatorname{Hom}_{\operatorname{\mathscr{D}M^{(0)}}(k)}(\operatorname{\hat{M}}(E),\mathbbm{1}\{n\})\simeq\operatorname{H}_{\mathrm{M}}^{n,n}(E)\simeq K_{n}^{M}(E).$
    Thus, one can associate to any symbol $\sigma=\{u_{1},\ldots,u_{n}\}\in K_{n}^{M}(E)$ a canonical map $\mu(\sigma):\operatorname{\hat{M}}(E)\rightarrow\mathbbm{1}\{n\}$.
    Using the diagonal map $\delta:\operatorname{Spec}(E)\rightarrow\operatorname{Spec}(E)\times_{k}\operatorname{Spec}(E)$, or equivalently, the multiplication map $E\otimes_{k}E\rightarrow E$, one obtains the required morphism of generic motives:
    $\gamma_{\sigma}:\operatorname{\hat{M}}(E)\xrightarrow{\delta_{*}}\operatorname{\hat{M}}(E)\otimes\operatorname{\hat{M}}(E)\xrightarrow{\mathrm{Id}\otimes\mu(\sigma)}\operatorname{\hat{M}}(E)\{n\}.$
@@
    Residues.– Let $(E,v)$ be a valued function field over $k$. Let $\mathcal{O}_{v}$ and $\kappa(v)$ be the associated ring of integers and residue field respectively. This induces a closed immersion between essentially smooth $k$-schemes $i:\mathcal{Z}=\operatorname{Spec}(\kappa(v))\rightarrow\operatorname{Spec}(\mathcal{O}_{v})=\mathcal{X}$, with open complement $j:\mathcal{U}=\operatorname{Spec}(E)\rightarrow\operatorname{Spec}(\mathcal{O}_{v})$. The Gysin triangle extends to this general setting and yields a (split) homotopy exact sequence in the stable $\infty$-category $\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M^{eff}_{gm}}(k)$:
@@
    The map $\partial_{v}$ is referred to as the residue map associated with $(E,v)$.
This concludes the description of the functor $\rho$ of (3.3.1.a) on morphisms. We refer the reader to [Dég08b] for the detailed proof of the relations.
3.3.5.
Let $\mathbf{E}$ be a motivic complex in $\operatorname{\mathscr{D}M}(k)$. Seen as a constant pro-object, it represents a functor $\operatorname{\hat{h}}_{\mathbf{E}}:\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M}(k)}(-,\mathbf{E})$, that can be restricted to the category of generic motives. Composing further with the functor $\rho$, one obtains a functor:
$\hat{\mathbf{E}}:\tilde{\mathcal{E}}_{k}\xrightarrow{\rho^{\mathrm{op}}}\operatorname{\mathscr{D}M^{(0)}}(k)^{\mathrm{op}}\xrightarrow{\operatorname{\hat{h}}_{\mathbf{E}}}\mathrm{A}b$
which is, by definition, a cycle premodule. We call it the Rost transform of $\mathbf{E}$. One can check that it is in fact a cycle module. Furthermore, according to [Dég11, Th. 3.7], one gets:
@@
$\mathcal{R}:\operatorname{HM}(k)=\operatorname{\mathscr{D}M}(k)^{\heartsuit}\rightarrow\operatorname{\mathscr{M}Cycl}_{k},\mathbf{E}\mapsto\hat{\mathbf{E}},$
where $\operatorname{\mathscr{M}Cycl}_{k}$ is Rost’s category of cycle modules, is an equivalence of categories, with a quasi-inverse given by the functor that associates to a cycle module $M$ its 00-th Chow group functor $\underline{H}M:=A^{0}(-,M)$.
@@
Note the last assertion was not proved in loc. cit. but it easily follows since, after inverting $p$, both categories $\operatorname{HM}(k)[1/p]$ and $\operatorname{\mathscr{M}Cycl}_{k}[1/p]$ become invariant under purely inseparable extensions of the base field $k$.
@@
According to 3.1.7, for any motivic complexe $\mathbf{E}$, one obtains:
$\hat{\mathbf{E}}(K,n)=\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M}(k)}(\operatorname{\hat{M}}(K)\{n\},\mathbf{E})\simeq\operatorname{\underline{H}}_{0,-n}(\mathbf{E})(K).$
In particular, the functor $\mathbf{E}\mapsto\hat{\mathbf{E}}$ factors through the homological functor $\operatorname{\underline{H}}_{0,*}$ associated with the homotopy $t$-structure on $\operatorname{\mathscr{D}M}(k)$.
Moreover, the abelian group $\hat{\mathbf{E}}(K,n)$ can be interpreted as the fiber of the Nisnevich sheaf $\operatorname{\underline{H}}_{0,-n}(\mathbf{E})$ at the point (in the topos-theoretic sense) of the Nisnevich site $\mathrm{Sm}_{k}$ corresponding to the function field $K$. As shown in 3.1.10, generic motives can be viewed as points of the $\infty$-$t$-category $\operatorname{\mathscr{D}M}(k)$, equipped with its homotopy $t$-structure. The Rost transform functor $\mathcal{R}$ can thus be interpreted as the restriction of homotopy modules to this category of points. Remarkably, restricting a homotopy module $M$ to this enriched category — via the functoriality encoded by $\rho$ — is sufficient to fully reconstruct $M$, as well as morphisms of homotopy modules.
[…]
4. Computations of generic motives
Throughout this section, we adopt the notation for motivic complexes and generic motives as recalled in the previous two sections. Most notably, we make use of the functoriality of generic motives as described by the functor (3.3.1.a).
4.1. Initial remarks about generic motives
4.1.1.
In [Dég02, §9.3.2], the question was raised whether morphisms of type (D*) generate all morphisms between generic motives. More precisely, is the functor $\rho$, defined in the previous section (3.3.1.a), full and/or faithful ?
This question has a negative answer (see 4.2.6), but there are still interesting cases where the functor $\rho$ does induce an isomorphism on morphisms.